# Bilinear forms with trace functions

### v2

@@ 1.1. Bilinear sums of trace functions
This is the same context as in various previous papers (see [FI] by Friedlander and Iwaniec, [FMAnn] by Fouvry and Michel and [KMSAnn], [Pisa] by Kowalski, Michel and Sawin). The key achievement of the present work is to obtain estimates as good as those of these papers, but in much greater generality. More precisely, whereas it was previously always assumed that $K$ was of a specific type (essentially some exponential sum of specific type, such as hyper-Kloosterman sums), we are now able to succeed with qualitative conditions on $K$, and these conditions are, on the one hand, known to hold for a very wide variety of trace functions and, on the other hand, extremely robust under a variety of transformations.
@@ 1.4. Cubic toroidal moments
It is also useful to decompose $M_{a,b,c}(q)$ along the even and odd characters, so we write
$M_{a,b,c}(q)=M^{+1}_{a,b,c}(q)+M^{-1}_{a,b,c}(q)$
where
$M^{\varepsilon}_{a,b,c}(q)=\frac{1}{q-1}\sum_{\stackrel{{\scriptstyle{\scriptstyle\chi\,(\mathrm{mod}\,{q})}}}{{{\scriptstyle\chi(-1)=\varepsilon}}}}L(1/2,\chi^{a})L(1/2,\chi^{b})L(1/2,\chi^{c}),\quad\quad\varepsilon=+1,\ -1.$
@@
There exists $\eta>0$ and there exist polynomials $P_{a,b,c}=1$ and $P_{a,b,-c}$ such that
$\deg(P_{a,b,-c})$ $=\delta_{a,c}=\begin{cases}1&\text{ if }a=c,\\ 0&\text{ if }a\not=c,\end{cases}$
$P_{a,b,-c}$ $\text{ has positive leading term},$
such that the estimates
$M^{+1}_{a,b,\varepsilon c}(q)$ $\geqslant P_{a,b,\varepsilon c}(\log q)+O(q^{-\eta}),$
$M_{a,a,\varepsilon c}(q)$ $=P_{a,a,\varepsilon c}(\log q)+O(q^{-\eta}),$
@@
In order for this result to be applicable, we will also prove that almost all sheaves $\mathscr{K}_{a,b,c}$ are either gallant or oxozonic, using the work of Katz.
@@
$|\{\chi\,(\mathrm{mod}\,{q}),\ \chi\not=1,\ L(1/2,\chi^{a})L(1/2,\chi^{b}),\ \varepsilon(\chi)\in I\}|\geqslant c(q-2),$
@@ 3.1. Goursat’s Lemma
We recall a version of the classical lemma of Goursat.
@@
$H_{1}\times H_{2}\subset G\subset\{(x,y)\in G_{1}\times G_{2}\,\mid\,\varphi(xH_{1})=yH_{2}\}.$
@@
The kernel of the restriction of $p_{1}$ (resp. $p_{2}$) to $G$ is identified with a subgroup $H_{2}$ of $G_{2}$ (resp. a subgroup $H_{1}$ of $G_{1}$). One then checks that the image of the natural morphism
@@ 3.2. A criterion for vanishing of coinvariants
(since either $\varrho_{j}(x^{\prime})$ or $\varrho_{j}(y^{\prime})$ equals $1$); this completes the induction (compare with the proof of [31, Lemma 3.3], due to Serre).
@@ 3.4. A property of gallant sheaves
Let $W=\overline{\mathbf{Q}}_{\ell}[\operatorname{Aut}(S)]$ be the regular representation of $\operatorname{Aut}(S)$. Let $\mathscr{H}$ be the sheaf on $\mathbf{A}^{1}_{k}$ associated to $W$. Since the monodromy group of $\mathscr{H}$ is finite, this sheaf geometrically semisimple. After replacing $k$ by a finite extension, if necessary, we may assume that $\mathscr{H}$ splits as a direct sum of geometrically irreducible sheaves. Let then $I$ be the finite set of geometrically irreducible components which arise in this decomposition.
@@ 3.6. Diagonal estimates
We will use repeatedly the following simple combinatorial lemma. By “graph”, we mean undirected graphs without multiple edges, but where loops are allowed.
@@ 9.5. Tannakian operations
for $a\in\mathbf{F}_{q}^{\times}$. It would be interesting to see if such examples have concrete analytic applications.
@@ 9.6. Hypergeometric sheaves
Nevertheless, we discuss briefly the exceptional cases $r\in\{4,8,9\}$ since these may occur naturally. From the combination of several results of Katz [ESDE], namely Theorem 8.8.1, Theorem 8.8.2, Theorem 8.11.2 and Corollary 8.11.2.1, Theorem 8.11.3, Theorem 10.1.3, Theorem 10.7.1, Theorem 10.8.1 and Theorem 10.9.1 from loc. cit., one obtains the following:
@@
  • $r=4$, there exists a character $\eta$ such that the multisets $\eta\cdot\text{${\chi}$}$ and $\eta\cdot\text{${\varrho}$}$ are both invariant under the inversion map $\xi\mapsto\xi^{-1}$ and
@@
    where $\chi_{1/2}$ is the unique character of exact order $2$. In that case we have $G^{0}=\mathbf{SO}_{4}$ acting on $\mathrm{St}_{4}$ and the geometric monodromy group $G_{\eta}$ of the twisted sheaf $\mathscr{F}\otimes\mathscr{L}_{\eta}$ equals and
    $G_{\eta}=\begin{cases}\mathbf{SO}_{4}&\hbox{ if }\eta^{r}\prod_{i=1}^{r}\chi_{i}=1,\\ \mathbf{O}_{4}&\hbox{ if }\eta^{r}\prod_{i=1}^{r}\chi_{i}=\chi_{1/2}.\end{cases}$
  • $(r,t)=(8,2)$, $|k|\equiv 1\,(\mathrm{mod}\,{4})$ and there exists characters of $k^{\times}$}, $\eta$, $\chi$ with $\chi$ of not exact order $4$, such that
    $\eta\cdot\text{${\chi}$}=\{\chi,\overline{\chi}\}\cup\{\hbox{$3$-rd roots of }\chi,\overline{\chi}\}\,,\ \eta\cdot\text{${\varrho}$}=\{\chi_{1/4},\chi_{3/4}\},$
@@
    In that case one has $G^{0}=\mathbf{SL}_{2}\times\mathbf{SL}_{2}\times\mathbf{SL}_{2}$ acting on ${\mathrm{Std}_{2}}^{\otimes 3}$.
@@
    $\eta\cdot\text{${\chi}$}=\{\chi,\varrho,\xi\}\cup\{\hbox{$2$-nd roots of }\overline{\chi},\overline{\varrho},\overline{\xi}\},\ \eta\cdot\text{${\varrho}$}=\{1,\chi_{1/3},\chi_{2/3}\}$
@@
In the $r=t$ case, the monodromy group could be finite. Under the assumption that it is infinite, the results of Katz give the following result.
@@
$\prod_{i=1}^{r}\chi_{i}\times\prod_{j=1}^{t}\varrho_{j}^{-1}=\chi_{1/2}$
@@
$G_{\eta}=\begin{cases}\mathbf{SO}_{4}&\hbox{ if }\eta^{r}\prod_{i=1}^{r}\chi_{i}=1,\\ \mathbf{O}_{4}&\hbox{ if }\eta^{r}\prod_{i=1}^{r}\chi_{i}=\chi_{1/2}.\end{cases}$

### v3

@@ 1.1. Bilinear sums of trace functions
This is the same context as in various previous papers (see [FI] by Friedlander and Iwaniec, [FMAnn] by Fouvry and Michel and [KMSAnn], [Pisa] by Kowalski, Michel and Sawin). The key achievement of the present work is to obtain estimates as good as those of these papers, but in much greater generality. More precisely, whereas it was previously always assumed that $K$ was of a specific type (essentially some exponential sum of specific type, such as hyper-Kloosterman sums), we are now able to succeed with qualitative conditions on $K$, and these conditions are, on the one hand, known to hold for a wide variety of trace functions and, on the other hand, extremely robust under a variety of natural transformations.
@@ 1.4. Cubic toroidal moments
There exists $\eta=\eta(a,b,c)>0$ and $D_{a,b,\varepsilon c}\geqslant 1$ such that the estimates
$M_{a,b,\varepsilon c}(q)$ $\geqslant D_{a,b,\varepsilon c}+O(q^{-\eta}),$
$M_{a,a,\varepsilon c}(q)$ $=D_{a,a,\varepsilon c}+O(q^{-\eta}),$
@@
In order for this result to be applicable, we will also prove that almost all sheaves $\mathscr{K}_{a,b,c}$ are either gallant or oxozonic, using the works of Katz and Katz–Tiep.
@@
$|\{\chi\,(\mathrm{mod}\,{q}),\ \chi\not=1,\ L(1/2,\chi^{a})L(1/2,\chi^{b})\not=0,\ \varepsilon(\chi)\in I\}|\geqslant c(q-2),$
@@ 3.1. Goursat’s Lemma
We recall a version of the classical lemma of Goursat (see also the version by Serre in [SerreBook, Prop. 1.6]).
@@
$H_{1}\times H_{2}\subset G=\{(x,y)\in G_{1}\times G_{2}\,\mid\,\varphi(xH_{1})=yH_{2}\}.$
@@
[…]
@@ 3.2. A criterion for vanishing of coinvariants
(since either $\varrho_{j}(x^{\prime})$ or $\varrho_{j}(y^{\prime})$ equals $1$); this completes the induction (compare with the proof of [32, Lemma 3.3], due to Serre).
@@ 3.4. A property of gallant sheaves
Let $W=\overline{\mathbf{Q}}_{\ell}[\operatorname{Aut}(S)]$ be the regular representation of $\operatorname{Aut}(S)$. Let $\mathscr{H}$ be the sheaf on $\mathbf{A}^{1}_{k}$ associated to $W$. Since the monodromy group of $\mathscr{H}$ is finite, this sheaf is geometrically semisimple. After replacing $k$ by a finite extension, if necessary, we may assume that $\mathscr{H}$ splits as a direct sum of geometrically irreducible sheaves. Let then $I$ be the finite set of geometrically irreducible components which arise in this decomposition.
@@ 3.6. Diagonal estimates
We will use repeatedly the following simple combinatorial lemma. By “graph”, we mean an undirected graph without multiple edges, but where loops are allowed.
@@ 9.2. Birationality
9.3. Twists
If $\mathscr{F}$ is gallant, and if $\mathscr{L}$ is a constructible $\ell$-adic sheaf of generic rank $1$ on $\mathbf{A}^{1}$ over $k$, lisse on an open dense subset $U$, then the twisted sheaf $\mathscr{L}\otimes\mathscr{F}$, which has trace function
$x\mapsto t_{\mathscr{L}}(x;k)t_{\mathscr{F}}(x;k)$
is also gallant.
This fact is immediate from the definition when $G^{0}$ is non-trivial, because the geometric monodromy group of $\mathscr{L}\otimes\mathscr{F}$ is the same as that of $\mathscr{F}$. If $G^{0}$ is trivial, we can argue as follows. Let $N$ be the core subgroup of $G$, and $H$ the preimage of $N$ in the Galois group of $k(T)$ for the homomorphism corresponding to $\mathscr{F}$. We then have, by restriction, a surjective homomorphism $\varrho\colon H\to N$ corresponding to $\mathscr{F}$ and a character $\chi\colon H\to\overline{\mathbf{Q}}_{\ell}^{\times}$ corresponding to $\mathscr{L}$. The subgroup $\varrho(\ker(\chi))$ is a normal subgroup of $N$ since $\varrho$ is surjective, so it is either central or equal to $N$, since $N$ is quasisimple (see [isaacs, Lemma 9.2]). If it were central, we would obtain a surjective homomorphism
$H/\ker(\chi)\to H/\varrho^{-1}(Z(N))\stackrel{{\scriptstyle\varrho}}{{\longrightarrow}}N/Z(N),$
which is impossible since $H/\ker(\chi)$ is abelian and $N/Z(N)$ is a non-abelian simple group.
So we deduce that $\varrho(\ker(\chi))=N$, and this implies that the image of the Galois representation associated to $\mathscr{L}\otimes\mathscr{F}$ contains $N$, so that this sheaf is gallant.
This means, for instance, that if we can apply our results to a function $K$ on $\mathbf{F}_{q}$, then they are also applicable to functions like
$K(x)\chi(g(x))e(f(x)/q)$
for any multiplicative character $\chi$ modulo $q$, and any rational functions $f$ and $g$ modulo $q$ which can be written as ratios of polynomials with degree uniformly bounded as $q$ varies (the last requirement ensuring that the complexity of the corresponding twisted sheaf remains bounded).
@@ 9.6. Tannakian operations
for $a\in\mathbf{F}_{q}^{\times}$. See [MiInv, FMAnnals, XiInv, XiIMRN] for situations where the problem of bounding such kind of bilinear sums might be useful.
@@ 9.7. Hypergeometric sheaves
Nevertheless, we discuss briefly the exceptional cases $r\in\{4,8,9\}$ since these may occur naturally. From the combination of several results of Katz [ESDE], namely Theorem 8.8.1, Theorem 8.8.2, Theorem 8.11.2 and Corollary 8.11.2.1, Theorem 8.11.3, Lemma 8.11.6, Theorem 10.8.1 and Theorem 10.9.1 from loc. cit., one obtains the following:
@@
  • $r=4$, and there exists a character $\eta$ such that the multisets $\eta\cdot\text{${\chi}$}$ and $\eta\cdot\text{${\varrho}$}$ are both invariant under the inversion map $\xi\mapsto\xi^{-1}$ and
@@
    where $\chi_{1/2}$ is the unique character of exact order $2$. In that case we have $G^{0}=\mathbf{SO}_{4}$ acting by the standard representation $\mathrm{St}_{4}$. The geometric monodromy group $G_{\eta}$ of the twisted sheaf $\mathscr{F}\otimes\mathscr{L}_{\eta}$ is given by the formula
    $G_{\eta}=\begin{cases}\mathbf{SO}_{4}&\hbox{ if }\prod_{i=1}^{4}\eta\cdot\chi_{i}=1\\ \mathbf{O}_{4}&\hbox{ if }\prod_{i=1}^{4}\eta\cdot\chi_{i}=\chi_{1/2},\end{cases}$
    where $\chi_{1/2}$ is the unique character of order $2$.
  • $(r,t)=(8,2)$ and one of the two holds
    1. (1)
      $|k|\equiv 1\,(\mathrm{mod}\,{4})$ and there exists characters of $k^{\times}$}, $\eta$, $\chi$ with $\chi$ not of exact order $4$, such that
      (9.3) $\eta\cdot\text{${\chi}$}=\{\chi,\overline{\chi}\}\cup\{\hbox{$3$-rd roots of }\chi,\overline{\chi}\}\,,\ \eta\cdot\text{${\varrho}$}=\{\chi_{1/4},\chi_{3/4}\},$
@@
    2. (2)
      $|k|\equiv 1\,(\mathrm{mod}\,{3})$ and there exists characters of $k^{\times}$}, $\eta$, $\chi$ with $\chi$ not of exact order $4$, such that
      (9.4) $\eta\cdot\text{${\chi}$}=\{\chi,\overline{\chi}\}\cup\{\hbox{$3$-rd roots of }\chi,\overline{\chi}\}\,,\ \eta\cdot\text{${\varrho}$}=\{\chi_{1/3},\chi_{2/3}\},$
      where $\chi_{1/3},\chi_{1/3}$ denote the two characters of $k^{\times}$} of exact order $3$.
    In these cases one has $G^{0}=\mathbf{SL}_{2}\times\mathbf{SL}_{2}\times\mathbf{SL}_{2}$ acting on ${\mathrm{Std}_{2}}^{\otimes 3}$.
@@
    (9.5) $\eta\cdot\text{${\chi}$}=\{\chi,\varrho,\xi\}\cup\{\hbox{$2$-nd roots of }\overline{\chi},\overline{\varrho},\overline{\xi}\},\ \eta\cdot\text{${\varrho}$}=\{1,\chi_{1/3},\chi_{2/3}\}$
@@
Proof.
Suppose $r=4>t$. By [ESDE]*Thm 8.11.3 (1) (2), we have $G^{0}=G^{0,der}$, and the sheaf is gallant unless rtr-t is even.
Suppose rtr-t is even. Then, by [ESDE]*Thm 8.11.3 (3), we have $G^{0}=\mathbf{SL}_{4},\ \mathbf{SO}_{4}$ or $\mathbf{Sp}_{4}$, depending on whether or not some $\mathscr{F}$ twisted by some $\mathscr{L}_{\eta}$ is self-dual or not, and on the type (symmetric or alternating) of the duality.
To be precise, by [ESDE]*Thm 8.8.1, a twist $\mathscr{L}_{\eta}\otimes\mathscr{F}$ is self-dual if and only if $\eta\cdot\text{${\chi}$}$ and $\eta\cdot\text{${\varrho}$}$ are both invariant under the inversion map $\xi\mapsto\xi^{-1}$. If this is not the case, then $G^{0}=\mathbf{SL}_{4}$. Otherwise, we have
$\eta^{r-t}\prod_{i=1}^{r}\chi_{i}\times\prod_{j=1}^{t}\varrho_{j}^{-1}=1\text{ or }\chi_{1/2},$
and by [ESDE]*Thm 8.8.2, the duality is alternating in the first case (and then $G^{0}=\mathbf{Sp}_{4}$) and symmetric in the second case (and then $G^{0}=\mathbf{SO}_{4}$).
Suppose we are in the symmetric case. The determinant of $\mathscr{F}\otimes\mathscr{L}_{\eta}$ is given by [ESDE]*Lemma 8.11.6; since $r-t\geqslant 2$, it is equal to $\mathscr{L}_{\Lambda_{\eta}}$ where
$\Lambda_{\eta}=\eta^{4}\prod_{i=1}^{4}\chi_{i}.$
This character is either $1$ or $\chi_{1/2}$, and the geometric monodromy group of $\mathscr{F}\otimes\mathscr{L}_{\eta}$ is correspondingly either $\mathbf{SO}_{4}$ or $\mathbf{O}_{4}$.
If $r=8$ or $9$, then by [ESDE]*Thm 8.11.3 (1) (2) (3) $G^{0}=G^{0,der}$ and the sheaf is gallant unless rt=6r-t=6.
If $(r,t)=(8,2)$, then by [ESDE]*Thm 8.11.3 (1) (2) (3) and [ESDE]*Theorem 10.8.1 the sheaf is gallant unless $(\eta.\text{${\chi}$},\eta.\text{${\varrho}$})$ is of the shape (9.3) or (9.4).
If $(r,t)=(9,3)$, then by [ESDE]*Thm 8.11.3 (1) (2) (3) and [ESDE]*Theorem 10.9.1 the sheaf is gallant unless $(\eta.\text{${\chi}$},\eta.\text{${\varrho}$})$ is of the shape (9.5). ∎
@@
In the $r=t$ case, the monodromy group could be finite. Under the assumption that it is infinite, the work of Katz provides the following:
@@
$\prod_{i=1}^{4}\chi_{i}\times\prod_{j=1}^{4}\varrho_{j}^{-1}=\chi_{1/2}$
@@
$G_{\eta}=\begin{cases}\mathbf{SO}_{4}&\hbox{ if }\prod_{i=1}^{4}\eta.\chi_{i}=1,\\ \mathbf{O}_{4}&\hbox{ if }\prod_{i=1}^{4}\eta.\chi_{i}=\chi_{1/2}.\end{cases}$